An angle measure (x+13) another angle complementary to it measure (2x+23) what's the value of x

Found 2 solutions by jim_thompson5910, mananth:
Answer by jim_thompson5910(35256)

An angle measure (x+13) another angle complementary to it measure (2x+23) what's the value of x


-------------------------------------------------------


Complementary angles add to 90, so,


(x+13)+(2x+23) = 90


x+13+2x+23 = 90


3x+36 = 90


3x = 90-36


3x = 54


x = 54/3


x = 18

Answer by mananth(16946)

The sum of complimentary angles is 90 degrees
(x+13) another angle complementary to it measure (2x+23)
(x+13)+(2x+23)=90
3x+36=90
-36 -36
3x=54
/3
x=18